OpenStudy (anonymous):

In triangle RST, RS=10, RT=15, and measurement R=32. In triangle UVW, Uv==12, UW=18, and measurement U=32. Are the polygons similar? If they are, write a similarity statement and give the similarity ratio?

OpenStudy (anonymous):

\[\frac{RT}{RS} : \frac{UW}{UV}\] \[\frac{15}{10}:\frac{18}{12}\] simplify both and gt th ratio...do u knw to get the ratio

OpenStudy (anonymous):

Yes Thanks !